OpenStudy (anonymous):

You are saving $5 each week to attend a summer program that costs $300. You already have $84. You have also received a $50 scholarship for the program. How many weeks of saving do you need?

OpenStudy (jebonna):

300 - 84 = 216 216 - 50 = 166 166/5 = 33.2 Around 33 weeks of saving is needed

OpenStudy (anonymous):

34 weeks

OpenStudy (anonymous):

as you would still not have enough money at 33 weeks of saving so you would need to save for 34 weeks

OpenStudy (jebonna):

Yaa